5. Evaluate f, where f(x + iy) = x2 + iya and y is the line joining 1 to i.
Answer #1

Let me discuss this problem part by part

5 I= S + f(xtiy) = 1²+iy2 and si the line joining 10 8 11 I to u 12 so T = f f 2 ul., (1,0) to (0, 1) Equation of line joinin

0 = (1-ig (x2+x-i-zxi +i) dx Ő 10 3 = (1-i) + -i +ix) 11 12 = (1-1) [0 - 5 - 1 +9=*] = (1-1) = (1+i) Ś (1-i) (1+i) - 1/3 (1-(